Suppose a uniform random variable can be used to describe the outcome of an experiment with the outcomes ranging from

20 to 70. What is the probability that this experiment results in an outcome less than 30​?
Answer #1

Solution :

Given that,

a = 20

b = 70

P(x < c) = (c - a) / (b - a)

P(x < 30) = (30 - 20) / (70 - 20) = 0.2

Probability = 0.2